### v4

# Faster and Memory-Efficient Training of Sequential Recommendation Models for Large Catalogs

Faster and Memory-Efficient Training of Sequential Recommendation Models for Large Catalogs
Maxim Zhelnin zhelninmax@gmail.com MWS AIMoscowRussia , Dmitry Redko dmitryredko444@gmail.com Applied AIMoscowRussia , Daniil Volkov danvol121@gmail.com Applied AIMoscowRussia , Anna Volodkevich volodkanna@yandex.ru Sber AI Lab, Applied AIMoscowRussia , Petr Sokerin sokerinpo@gmail.com Applied AIMoscowRussia , Valeriy Shevchenko escape756@gmail.com IVIMoscowRussia , Egor Shvetsov e.shvetsov@applied-ai.ru Applied AIMoscowRussia , Alexey Vasilev alexxl.vasilev@yandex.ru Sber AI LabMoscowRussia , Darya Denisova duny.explorer@gmail.com Sber AI LabMoscowRussia , Ruslan Izmailov lord.rik@yandex.ru SberMoscowRussia and Alexey Zaytsev likzet@gmail.com Applied AIMoscowRussia
Sequential Recommendation, Cross-Entropy Loss Optimization, GPU Memory Constraints, Negative Sampling, Training Efficiency, Softmax Saturation, Memory-Efficient Training
$c$s: Information systems Recommender systems
Sequential recommendations (SR) with transformer-based architectures are widely adopted, and SR models require frequent retraining to adapt to changing user preferences. However, training transformer-based SR models incurs high computational cost from scoring large item catalogs (often thousands of items). This is mainly due to cross-entropy loss, where peak memory scales with catalog size, batch size, and sequence length. To reduce memory, practitioners combine cross-entropy (CE) with negative sampling, lowering explicit memory needs of the final layer. Yet few negative samples degrade performance; increasing negatives and batch size improves results but quickly exceeds industrial GPU memory ( 40Gb).
In this work, we introduce $\mathrm{CCE}^{-}$, a GPU-efficient implementation of CE with negative sampling. $\mathrm{CCE}^{-}$ accelerates training by up to 2× and reduces memory by more than 10×. The saved memory enables higher accuracy on datasets with large catalogs versus models trained with standard PyTorch losses. We also analyze key memory-related hyperparameters and show the need to balance them: scaling both the number of negative samples and batch size yields better results than maximizing only one. To support adoption, we release a Triton kernel implementing the method efficiently. [^1] Code, reproducibility materials, and all scripts for generating figures are available at https://github.com/On-Point-RND/MemoryEfficientSRS
1. Introduction
User actions can be modeled sequentially, conditioning on past actions in diverse scenarios, ranging from a mechanical series of actions in a banking app (boka2024survey) to a mood- and neighborhood-influenced selection of music tracks (krause2014music). Consequently, Sequential Recommender Systems (SRS), which model user interactions as sequences, have become essential components of modern recommendation pipelines. SRS effectively incorporates ideas from natural language processing (NLP) through attention-based architectures such as SASRec (kang2018self) and BERT4Rec (sun2019bert4rec), as well as the cross-entropy loss function (klenitskiy2024does; sun2019bert4rec; mezentsev2024scalable).
The deployment of SRS often faces two challenges: (1) high memory demands due to large catalogs, (2) a need for frequent retraining to acquire recent user actions. Although peak memory usage was extensively addressed in the literature, training speed did not receive as much attention. In this paper, we consider both problems.
Peak memory
Peak memory is one of the main computational constraints when cross-entropy loss (CE) is applied to large catalogs. The memory footprint at the final layer for logits can be estimated as $bs\cdot sl\cdot|V|$ [^1] To illustrate, consider a batch of 256 sequences, each of length 100, with a catalog containing $10^{6}$ items. The resulting logit tensor would require the storage of $256\cdot 100\cdot 10^{6}$ elements, equating to approximately 100 GB of memory., where $bs$ is the batch size, $sl$ sequence length and $|V|$ is the catalog or vocabulary size, typically the most significant contributor. This final layer uses up to 90% of the combined memory footprint, gradients, and activations, when training modern Large Language Models (LLMs) (wijmans2024cut). However, in SRS, $|V|$ can reach millions of items (zivic2024scaling), surpassing even LLMs, which have vocabulary sizes orders of magnitude smaller: 32K, 152K, and 256K for Llama2-7B (touvron2023llama), QWEN (bai2023qwen), and Gemma-7B (team2403gemma), respectively.
Figure 1. Here, we compare several methods CE, $\mathrm{CE}^{-}$, CCE, $\mathrm{CCE}^{-}$, and SCE in terms of (1) NDCG@10, (2) Training time per epoch in seconds, and (3) Memory consumption in Gb for the SASRec model across six datasets. We demonstrate the best performance achieved through optimized hyperparameters. We can see that CE, $\mathrm{CE}^{-}$ require much more time and memory usage.
Table 1. NDCG@10 metric values for the SASRec model trained using different loss functions. The best metric value is in bold, and the second-best is underlined.
Dataset BCE CE CE- CCE CCE-
Megamarket 0.0017 0.0052 0.0117 0.0181 0.0168
Zvuk 0.0046 0.0797 0.0675 0.0940 0.0852
30Music 0.0004 0.0891 0.0939 0.1393 0.1436
Gowalla 0.0197 0.0225 0.0278 0.0241 0.0279
Beauty 0.0094 0.0114 0.0123 0.0104 0.0116
ML-20M 0.0557 0.0510 0.0513 0.0511 0.0523
This memory footprint has been a key driver for early SRS adopting the binary cross-entropy loss (BCE) (kang2018self) and the cross-entropy loss with negative sampling ($\mathrm{CE}^{-}$(sun2019bert4rec; klenitskiy2023turning). $\mathrm{CE}^{-}$ directly addresses memory constraints caused by large item catalogs (klenitskiy2023turning; mezentsev2024scalable). Instead of scoring all items during training, a subset of negative samples ($ns$) is used, where $ns\ll|V|$, to approximate the full CE loss. This reduces computational complexity and memory requirements but introduces a hyperparameter trade-off between batch size, sequence length, and number of negative samples (ns) in memory-constrained scenarios. Prior work indicates that maximizing $ns$ in $\mathrm{CE}^{-}$ or using CE can improve performance of the models (di2024theoretical; klenitskiy2023turning; guo2024scaling). However, empirical results demonstrate that $\mathrm{CE}^{-}$ can sometimes outperform full CE for some datasets (mezentsev2024scalable; klenitskiy2023turning) and the performance of models trained with $\mathrm{CE}^{-}$ shows a non-linear dependence on nsns(prakash2024evaluating). This creates an open question of whether practitioners should maximize $ns$ given a fixed memory budget or prioritize other memory-related parameters. We will refer to this question as the first question Q1.
Training Speed
Although sampling strategies such as $\mathrm{CE}^{-}$ reduce memory consumption, they also improve training speed by requiring fewer operations, partially addressing the problem of frequent retraining SRS. However, there are other approaches for faster training: quantization (wortsman2023stable), activation sparsification (hu2024accelerating), and finally gradient sparsification (alistarh2018convergence). The latter one is usually overlooked, even though gradients usually require twice as many computations as the forward path. In our work, we demonstrate that gradient sparsification is especially suitable for SRS, provide a formal analysis in Section 3.2, and study how sparse gradients accelerate model training and their impact on accuracy degradation. We will refer to it as our second research question Q2.
Hardware-friendly Low-Level Efficient Implementations
Computational efficiency critically depends on the co-design of architecture, hardware, and low-level implementations. While hardware is often fixed, efficiency can be optimized through architectures better aligned with underlying hardware (shvetsov2024quantnas) or specialized low-level optimizations. Though highly specific, such optimizations yield significant computational benefits. Examples include FlashAttention (dao2022flashattention) for efficient attention computation and Liger kernels (hsu2024liger) for optimized CE computation. Among them, Cut-Cross Entropy (CCE) was recently developed for LLMs; it helps to alleviate memory bottlenecks during training when CEloss is employed. However, until now, the CCE has remained unexplored for SRS models. Furthermore, our analysis reveals the importance of negative sampling, which is not supported by the original CCE. Development and analysis of CCE with negative sampling is our third research question - Q3, more precisely: Can we design a CCE alternative that is faster and more memory efficient for SRS?
Our work addresses these questions through the following contributions:
  1. (1)
    Comprehensive Memory-Accuracy Trade-off Analysis: To address Q1, we conducted a systematic investigation into the relationships among accuracy of SASRec, batch size ($bs$), sequence length ($sl$), the number of negative sample size ($ns$) across six datasets with varying item catalog sizes ($|V|$). Our findings indicate that simply maximizing $ns$ in $\mathrm{CE}^{-}$ or employing CE to encompass the entire item catalog $|V|$ leads to suboptimal performance in memory-constrained environments. Instead, the choice between $\mathrm{CE}^{-}$ and CE requires a data set-specific calibration. We provide analysis for configuring these parameters to balance accuracy and memory efficiency.
    Furthermore, we make publicly available a dataset comprising 997 data points, each representing SASRec trained across the considered datasets. For every configuration, we log $bs$, slsl, nsns, memory consumption, and various metrics along the training steps. This dataset enables reproducible analysis of memory-accuracy trade-offs.
  2. (2)
    Sparse Gradients For training acceleration: To address Q2, we conducted an analysis of the classification layer and observed that the matrices involved in gradient computations contain numerous elements with low magnitude. Due to data format limitations [^2] The smallest value representable in FP16 is $5.96\cdot 10^{-8}$., this leads to high sparsity of the gradient matrices. Our experiments demonstrate that further gradient filtering preserves model quality while accelerating training. The mathematical analysis in Section 3.2 further confirms that this phenomenon should generalize across sequential recommendation (SRS) models trained with cross-entropy (CE) loss and large catalog sizes.
  3. (3)
    Development of Faster $\mathrm{CCE}^{-}$ kernel: We are the first to adapt and demonstrate that CCE is exceptionally well-suited for training SRS models when the CE loss is required. Furthermore, motivated by the findings of our first contribution on the significance of negative sampling for certain data sets, we developed $\mathrm{CCE}^{-}$, a GPU-optimized variant of the loss CE with negative sampling. We demonstrate that $\mathrm{CCE}^{-}$ outperforms CCE on 4/6 datasets and achieves up to 89% faster training as it is presented in Figure 1.
  4. (4)
    Memory-Efficient Metric Improvement: By reallocating the memory savings achieved through the use of CCE and $\mathrm{CCE}^{-}$ to increase batch size $bs$, sequence length $sl$, and the number of negative samples $ns$, we attain up to a 30% improvement in accuracy while maintaining the same computational budget.
2. Related work
Cross-Entropy and sampled CE for SRS. Sequential recommendation has seen significant advancements in recent years. One pivotal observation from  (klenitskiy2023turning) demonstrated that SASRec’s performance improves when trained with cross-entropy loss instead of binary cross-entropy. Unfortunately, computing cross-entropy over all items is computationally prohibitive. To address this, researchers adopted sampled cross-entropy loss, which uses a subset of negative items. This approach was theoretically justified in  (di2024theoretical), showing that increasing negative samples enhances performance. However, negative sampling has non-linear performance scaling. Despite its benefits, negative sampling amplifies popularity bias, as popular items are more likely to be sampled  (prakash2024evaluating). Consequently, model performance becomes highly dependent on dataset distribution and sampling strategy. Furthermore,  (prakash2024evaluating) highlights that performance does not scale linearly with the number of negative samples, which complicates the selection of hyperparameters. To address this problem, we perform a careful and extensive study of memory-dependent hyperparameter analysis for optimal accuracy-memory tradeoff (Section 5.1). Prior work on scaling for SRS models  (guo2024scaling; zhang2024scaling; zivic2024scaling; xu2025efficient) emphasizes the benefits of scaling sequence length and model depth. However, our experiments reveal that expanding a single factor [^3] Factors include batch size, sequence length, and number of negative samples. yields limited gains, underscoring the need for multi-dimensional scaling. Unlike  (mezentsev2024scalable), our work focuses only on feature-scaling [^4] Feature scaling refers to adjusting batch size, sequence length, etc. effects while keeping model architecture fixed.
Computationally efficient approaches. Advanced sampling techniques for balancing accuracy and memory were proposed in  (mezentsev2024scalable). Liger-Kernel  (hsu2024liger), implemented in the Triton language  (tillet2019triton), addresses memory consumption issues through the use of FusedLinearCrossEntropy (FLCE). While theoretical convergence guarantees for sparse gradient methods were established in  (alistarh2018convergence), it’s application for efficient training remains limited  (chekalina2024sparsegrad; wijmans2024cut). Building on FLCE memory-efficient cross-entropy with saturated gradient filtering beyond numerical precision was proposed (CCE) for LLMs  (wijmans2024cut). Since saturated gradients do not fit numerical precision and turn to zero, their removal does not affect model convergence. In this work, we analyze whether the removal of low-magnitude non-zero gradients would help to accelerate model training with minimal performance degradation.
Possible further directions and trends. Even when isolating experiments to feature scaling, there remains significant flexibility in training strategies. For example, work  (li2022stability) demonstrated that training GPT-like models with short sequences initially and gradually increasing them enables stable, efficient training with larger batch sizes and learning rates. Similarly, authors of  (di2024theoretical) found that starting with fewer negative samples and increasing their count later boosts SRS model performance. Moreover, based on our results, an adaptive threshold to prune gradients can be employed. While these adaptive strategies are promising, we focus on studying simpler non-adaptive approaches to establish a stronger foundation for future research.
3. Methods
Figure 2. Illustration of CE, CCE and $\mathrm{CCE}^{-}$ implementation. CE: Materializes full logits in GPU memory, matmul and softmax are computed separately $\rightarrow$ high GPU memory usage, full $C\times E$ matmul, extra data transfers between HBM and SRAM lead to higher time delay. CCE : No logits materialization, fused matmul + softmax $\rightarrow$ store only positive logits and LSE vector, reduced HBM $\leftarrow\rightarrow$ SRAM transfers, but still full CxE matmul. $\mathrm{CCE}^{-}$: No logits materialization, fused matmul + softmax on sampled items $\rightarrow$ store only positive logits and LSE vector, reduced HBM $\leftarrow\rightarrow$ SRAM transfers, matmul only for sampled columns of C and corresponding E.
3.1. Cut-Cross-Entropy for RecSys
Our proposed method, $\mathrm{CCE}^{-}$, builds upon CCE, which itself extends the Liger-Kernel (LG) (hsu2024liger). Below, we provide a brief review of both $\mathrm{CCE}^{-}$ and LG, highlighting their key differences. We then focus on how $\mathrm{CCE}^{-}$ introduces specific improvements tailored for sequential recommendation systems.
During training, we deal with two types of memory: GPU memory (High Bandwidth Memory, HBM) and on-chip shared memory (Static Random-Access Memory, SRAM). HBM is large but relatively slow and is used to store data for computations. GPU kernel loads input data from HBM into SRAM to perform computations. SRAM is significantly faster than HBM but has much less capacity. Therefore, only small chunks of input data can be loaded into SRAM at a time. Once computations are complete, the output is saved back to HBM. While there are more details, see (dao2022flashattention),  (wijmans2024cut), this structure shapes the main goal: reduce the amount of data loaded into GPU memory. It can be achieved by minimizing the overall amount of memory used by a model or optimizing transfers between HBM and SRAM.
The main idea behind LG is that it processes hidden states in smaller chunks, therefore reducing peak memory usage. It applies a linear classification head to each chunk, computes the logits, and then feeds these logits into a GPU to calculate partial losses with a special kernel. These losses are used to estimate the gradients for the weights of the classification head and the chunked hidden states. While chunking reduces memory usage, increasing the number of chunks introduces latency overhead.
CCE enhances LG through the fusion of logit and softmax computations within a single GPU kernel. By executing intermediate operations in the high-speed on-chip shared memory (SRAM), this approach only requires the materialization of logits for correct items and the log-sum-exp (LSE) vector. This eliminates the need to store the entire logit tensor in the global GPU memory, resulting in a substantial reduction in GPU memory usage during both forward and backward propagation. To implement this optimization, new Triton kernels were developed, effectively balancing computational performance and memory efficiency.
In this study, we incorporated CCE into sequential recommendation transformer models to evaluate its effectiveness for recommendation tasks. For this purpose, the three-dimensional output tensor of embeddings produced by the final transformer block is flattened into a matrix $\mathbf{E}\in\mathbb{R}^{N\times D}$, where $D$ represents the hidden size and $N$ corresponds to the total number of items, calculated as the product of the batch size $bs$ and the sequence length $sl$. The CE loss for the set of items of size $|V|$ is written as follows:
(1) $\mathcal{L}$ $=-\dfrac{1}{N}\sum_{i=1}^{N}\mathbf{l}_{i},$
(2) $\mathbf{l}$ $=(\mathbf{C}^{\mathrm{T}}\mathbf{E})_{\mathbf{x}}-\mathrm{log}\sum_{j=1}^{V}\exp{(\mathbf{C}^{\mathrm{T}}\mathbf{E})},$
where $\mathbf{C}\in\mathbb{R}^{D\times|V|}$ represents the linear classifier, projecting the embeddings $\mathbf{E}$ into the output space. The first term $(\mathbf{C}^{\mathrm{T}}\mathbf{E})_{\mathbf{x}}$ corresponds to the vector of logits for the correct items with indices $\mathbf{x}=(x_{1},\ldots,x_{N})$, which is computed as $(\mathbf{C}^{\mathrm{T}}\mathbf{E})_{\mathbf{x}}=[C_{x_{1}}^{\mathrm{T}}E_{1},\ldots,C_{x_{N}}^{\mathrm{T}}E_{N}]$, $C_{x_{j}}$ refers to the columns of matrix $\mathbf{C}$ associated with the correct items $x_{j}$, and $E_{j}$ denotes columns of the output embeddings matrix. The second term $\mathrm{log}\sum_{j=1}^{V}\exp{(\mathbf{C}^{\mathrm{T}}\mathbf{E})}$ represents a combination of the log-sum-exp operation and matrix multiplication.
Formulating the CE loss as Equation 2 allows it to be decomposed into calculating logits for correct items and performing a log-sum-exp operation on the full logits matrix $\mathbf{C}^{T}\mathbf{E}$. Consequently, all operations in the second term can be fused into a single GPU kernel, thereby avoiding the materialization of the large full logits matrix in the global GPU memory.
Besides, to accelerate computations during the backward pass, the Triton kernels for CCE are capable of performing gradient filtering at the final layer. This feature facilitates the efficient handling of sparsity in gradient matrices, which are common in recommender systems, as discussed in Section 3.2. Filtering is achieved by disregarding computations with values below a prescribed threshold. The default threshold is set to the numerical precision of the data format, although other values can also be used. Consequently, CCE not only reduces the memory footprint, but also accelerates training due to ignoring computations involving near-zero values.
Figure 3. An illustration of computing gradients for the weights of the final layer.
3.2. Sparse Gradients in Training with CE loss
In the recommender systems, the objective is to predict the next item from a predefined set $V$ of |V||V|{} items. Similar to the next-token prediction, there is exactly one correct class and |V|1|V|{}-1 incorrect classes. Thus, employing loss functions such as the CE loss provides a suitable estimate of model quality  (di2024theoretical).
Let us consider the gradient sparsity that arises in the final layer of a sequential recommendation model during training with the CE loss.
Let $\mathcal{L}=\mathcal{L}(\mathbf{p},\mathbf{t})$ denote the CE loss, that is used to measure discrepancy between model predictions $\mathbf{p}$ and target one-hot encoded labels $\mathbf{t}$. The predicted probability vector $\mathbf{p}$ is computed as:
$\mathbf{p}=\frac{\exp(\mathbf{C}^{T}E)}{\sum_{j=1}^{|V|}\exp(C^{T}_{j}E)},$
where $\mathbf{C}$ is the weight matrix of the final layer, $C_{j}$ refers to the columns of the matrix $\mathbf{C}$, and $E$ is the embedding vector.
For the incorrect label $y$, the gradient is $\frac{\partial\mathcal{L}}{\partial C_{yj}}=p_{y}e_{j}$, where $e_{j}$ is a component of the vector $E$, see Figure 3. When the item catalog size $|V|$ is large, the probability of incorrect labeling $p_{y}\rightarrow 0$ since the softmax function was applied, and so $\frac{\partial\mathcal{L}}{\partial C_{yj}}\rightarrow 0$. Because we have |V|1|V|-1 incorrect labels, most gradients remain near zero or saturate to zero based on data format, for example, for FP16 precision, the smallest positive number is $6\times 10^{-8}$.
Empirical Evaluation. To validate our analysis, we present histograms of gradient distributions before and after training on the Megamarket dataset. As shown in Figure 4, gradient values cluster near zero in both phases, confirming the persistence of sparsity throughout training.
Figure 4. Histograms of the gradient distributions at the final layer at the beginning (left) and end (right) of the training, the Megamarket dataset.
3.3. Cut-Cross-Entropy with Negative Sampling
As demonstrated in (klenitskiy2023turning), negative sampling offers dual advantages: it allows better accuracy on certain datasets and accelerates training compared to full vocabulary utilization. Motivated by this, we extend $\mathrm{CCE}^{-}$ by integrating negative sampling into it.
$\mathrm{CCE}^{-}$ Loss Formulation. The loss function operates on an embedding matrix $\mathbf{E}\in\mathbb{R}^{N\times D}$, which is obtained by flattening the 3D output tensor along its first two dimensions. During training, at each time step, the positive items corresponding to users are augmented with a dynamically sampled set of $ns$ negative items. This creates a compact item subset $V_{s}\subset V$ where: $|V_{s}|=1+ns$ (1 positive + $ns$ negatives) and $|V_{s}|\ll|V|$ (sampled subset size $\ll$ size of the full catalog of items ). The loss calculation is thereby confined to interactions within this compact subset, enabling efficient training, as discussed below. The $\mathrm{CCE}^{-}$loss is written as
(3) $\mathcal{L}_{NS}=-\dfrac{1}{N}\sum_{i=1}^{N}\mathbf{l}^{NS}_{i},$
(4) $\mathbf{l}^{NS}=(\mathbf{C}^{\mathrm{T}}\mathbf{E})_{\mathbf{x}}-\mathrm{log}\sum_{j=1}^{V_{s}}\exp{(\mathbf{C}^{\mathrm{T}}\mathbf{E})_{\mathbf{Inds}_{j}}}.$
The first term is identical to that in Eq. 2. The second term is a vector defined as
(5) $LSE=\mathrm{log}\sum_{j=1}^{V_{s}}\exp{(\mathbf{C}^{\mathrm{T}}\mathbf{E})_{\mathbf{Inds}_{j}}},$
which incorporates the log-sum-exp function and indexed matrix multiplication. This is expressed as $(\mathbf{C}^{\mathrm{T}}\mathbf{E})_{Inds_{j}}=[C_{x_{1j}}^{\mathrm{T}}E_{1},\ldots,\mathbf{C}_{x_{Nj}}E_{N}]$, where the set $Inds_{j}=(x_{1j},\ldots,x_{Nj})$ represents the indices of the columns of the matrix $\mathbf{C}$ used for negative sampling. Specifically, when $j=1$, the columns correspond to the correct items, while for $1<j<|V_{s}|+1$, the columns are related to negative examples associated with the embeddings $\mathbf{E}=[E_{1},\ldots,E_{N}]$ of the user items.
The correct logits are represented by the first term of Eq. 4 are materialized in global GPU memory and computed using an identical CCE kernel. The second term in the equation involves several operations: indexing $\mathbf{C}$ based on $\mathbf{Inds}_{j}$, performing the dot product and summing along the inner dimension for $\mathbf{C}_{x_{ij}}^{\mathrm{T}}E_{i}$, and computing the log-sum-exp operation.
To achieve a memory-efficient computation for this term, we developed a specialized kernel that fuses these operations. For this purpose, we adopt an approach for the online softmax computation. As a result, all intermediate operations are carried out on-chip using shared memory, while only the log-sum-exp vector is stored in the global memory (milakov2018onlinesoftmax; hsuliger).
During the backward pass, the gradients for the output embeddings and the matrices of the linear classifier are computed as follows:
(6) $\nabla E_{i}=\sum_{j=1}^{|V_{s}|}(\mathbf{S}\cdot\nabla LSE)_{ij}C_{x_{ij}},\quad\nabla C_{x_{ij}}^{\mathrm{T}}=(\mathbf{S}\cdot\nabla LSE)_{j}^{\mathrm{T}}\cdot E_{i},$
where softmax $\mathbf{S}=\mathrm{softmax}(\mathbf{C}^{\mathrm{T}}\mathbf{E})_{\mathbf{Inds}}$ and $\cdot$ denotes the row-by-row element-wise multiplication. As in the CCE, the matrices $\mathbf{S}$ and $\mathbf{S}\cdot\nabla LSE$ are not materialized in global memory; instead, most computations are performed in shared memory. We recompute $(\mathbf{C}^{\mathrm{T}}\mathbf{E})_{\mathbf{Inds}}$ at the backward pass and calculate the softmax $\mathbf{S}$ as $\mathbf{S}=\mathrm{softmax}(\mathbf{C}^{\mathrm{T}}\mathbf{E})_{\mathbf{Inds}}=\exp((\mathbf{C}^{\mathrm{T}}\mathbf{E})_{\mathbf{Inds}}-LSE)$. Only the gradient $\nabla LSE$ is retained in the global memory.
Algorithms 1 and 2 outline the computational process for the LSE vector from Equation 5 during the forward pass, the gradients from Equation 6 during the backward pass, and also clarify the data access scheme.
During the forward pass, the output embedding $\mathbf{E}$ is divided into the blocks $E_{n,D}$ of size $N_{B}\times D$. Since the inner dimension $D$ in transformer models for RecSys is typically not large (in  (kang2018self), $D<50$; in  (klenitskiy2023turning)(klenitskiy2024does), $D=64$; in  (guo2024scaling), $D<400$ ), division is not performed along the $D$ dimension. The matrix of the linear classifier $\mathbf{C}$ is loaded inside a for-loop by columns, determined by the block of indices $\mathbf{Inds}_{n,j}$ of size $N_{B}$.
During the backward pass, each kernel is executed for the blocks of the output embedding matrix $E_{n,D}$ of size $N_{B}\times D$ and the indices matrix $\mathbf{Inds}_{n,j}$ of size $N_{B}$. After the computations, the gradient $\nabla\mathbf{C}$ contains non-zero values only in columns corresponding to the indices provided by $\mathbf{Inds}$. Since $\mathrm{CCE}^{-}$ aims to achieve acceleration by restricting the catalog size through negative sampling, we do not implement gradient filtering for this method, unlike in CCE. However, gradient filtering can be integrated after Step 4 of Algorithm 2 in a similar manner to CCE.
Algorithm 1 Memory-efficient linear-log-sum-exp with negative sampling, forward pass
[…]
2:for blocks $\mathbf{E}_{n,D}$ do
3:  $\mathbf{m}=-\infty_{n}$, $\mathbf{d}=\mathbf{0}_{n}$ $\triangleright$ vectors of size $n$ in on-chip SRAM
4:  while $j<V_{s}$ do
5:   $\mathbf{c}_{n}=\mathbf{C}_{\mathbf{Inds}_{n,j},D}$ $\triangleright$ Indexed load of $\mathbf{C}$ into on-chip SRAM
6:   $\mathbf{o}_{n}=\sum_{d=1}^{D}\mathbf{E}_{n,d}\cdot\mathbf{C}_{\mathbf{x}_{n},d}$ $\triangleright$ Blockwise scalar product
7:   $\mathbf{m}_{new}=\max(\mathbf{m},\mathbf{o})$ $\triangleright$ Update the maximum value
8:   $\mathbf{d}=\mathbf{d}\cdot\exp(\mathbf{m}_{new}-\mathbf{m})+\exp(\mathbf{o}_{n}-\mathbf{m}_{new})$ $\triangleright$ Update the sum
9:   $\mathbf{m}=\mathbf{m}_{new}$ $\triangleright$ Reassign the maximum value
10:  end while
11:  $\text{LSE}_{n}=\mathbf{m}+\log\mathbf{d}$ $\triangleright$ Compute LSE values in on chip SRAM
12:end for
Algorithm 2 Memory-efficient linear-log-sum-exp with negative sampling, backward pass
1:for all blocks $\mathbf{E}_{n,D}$, $\mathbf{Inds}_{n,j}$ do
2:  $\mathbf{c}_{n}=\mathbf{C}_{\mathbf{Inds}_{n,j},D}$ $\triangleright$ Indexed load of $\mathbf{C}$ into on-chip SRAM
3:  $\mathbf{o}_{n}=\sum_{d=1}^{D}\mathbf{E}_{n,d}\cdot\mathbf{C}_{\mathbf{x}_{n},d}$ $\triangleright$ Blockwise scalar product
4:  $\mathbf{S}=\exp(\mathbf{o}_{n}-LSE_{n})$ $\triangleright$ Compute the softmax
5:  $\nabla E_{n,D}+=(\mathbf{S}\cdot\nabla LSE)_{nj}\cdot\mathbf{c}_{n}$ $\triangleright$ Update gradient for output embeddings
6:  $\nabla\mathbf{C}_{\mathbf{Inds}_{n,j},D}^{\top}+=(\mathbf{S}\cdot\nabla LSE)_{j}^{\mathrm{T}}\cdot E_{i}$ $\triangleright$ Indexed update gradient for the linear classifier
7:end for
4. Experimental Settings
4.1. Datasets
Table 2. Statistics for used datasets after preprocessing
Dataset Number ofaaaaaaaaaa Mean sequence
Items Users Interactions length
ML-20M 1515K 138138K 1218112181K $88.22$
Beauty 176176K 674674K 50285028K $7.46$
Gowalla 309309K 8585K 46614661K $54.98$
Zvuk 893893K 371371K 243412243412K $656.93$
30Music 909909K 4444K 2563925639K $577.71$
Megamarket 16611661K 350350K 114846114846K $327.88$
We use six datasets: 30Music  (turrin201530music), Zvuk  (shevchenko2024variability), MovieLens-20M (ML 20M)  (harper2015movielens), MegaMarket (shevchenko2024variability), Beauty  (mcauley2015image), and Gowalla  (cheng2013you). Our main focus is on datasets with a large item catalog: Zvuk, MegaMarket, Gowalla, and 30Music, which include more than 300K items. Moreover, Zvuk and MegaMarket are some of the largest open datasets in the recommendation domain  (shevchenko2024variability). MovieLens and Beauty are added because of their popularity, used in more than 14 and 45 papers according to (klenitskiy2024does). For Beauty, we employ its third version, which has the largest item catalog. The characteristics of all datasets after preprocessing are in Table 2.
4.2. Evaluation Methodology
Dataset Splitting:
Global Temporal Cutoff: To ensure alignment with real-world sequential prediction scenarios, we use a global temporal split with the last interaction as a target and ”by user” validation scheme (gusak2025time). We split interactions at the 90% timestamp quantile. All interactions before this cutoff form the training/validation pool, while interactions after constitute the test set. Validation Set: From the training pool, we randomly select 5% of users. For each selected user, their last interaction is withheld for validation. Test Set: For users in the post-cutoff test set, only their last interaction is retained for evaluation, while all previous ones are passed as model input.
Metrics:
Following prior work  (anelli2019discriminative; zhao2022revisiting; shevchenko2024variability), we prioritize NDCG@10 due to its strong correlation with overall ranking quality. Along with it, record other metrics, including Coverage and Surprisal (vasilev2024replay). All metrics are available in our publicly released dataset.
Models:
We primarily focus on the SASRec model, as it is more commonly used in practice and frequently outperforms Bert4Rec (klenitskiy2023turning; mezentsev2024scalable). However, to demonstrate the efficiency of CCE and $\mathrm{CCE}^{-}$ for other architectures, we also present the results of applying these loss functions to the training of BERT4Rec in Section 5.6
Baselines:
We analyze the efficiency of CCE and $\mathrm{CCE}^{-}$ for the SASRec by comparing metrics and computing resources expended with those for CE, $\mathrm{CE}^{-}$ in Section 5.2, for other sampling strategies in Section 5.4.1, and BCE in Section $5.$ Additionally, in Section 5.6 we analyze the efficiency of CCE and $\mathrm{CCE}^{-}$ for Bert4Rec.
4.3. Training Details
For all datasets, we trained models on a single Nvidia A100 GPU. Prior to our experiments, we tuned the number of self-attention blocks, attention heads, and hidden size. As a result, the model configuration was set to two self-attention blocks with two attention heads and a hidden size $D$ of $256$. The models were trained using the Adam optimizer with a learning rate of 1e31e-3. The training was performed with mixed-precision (float16) using the PyTorch Lightning trainer (micikevicius2017mixed).
5. Results
In the following, we present the results of our computational experiments and analyze them to address the questions Q1 and Q2 posed in Section 1 and evaluate the performance of CCE and $\mathrm{CCE}^{-}$. Specifically, (1) Section 5.1 contains the results of experiments carried out to study the effect of training hyperparameters on SASRec performance, where, among other insights, we demonstrate that sometimes increasing $ns$ does not lead to an improvement in the accuracy of the model; therefore, sometimes $\mathrm{CE}^{-}$ should be preferred over CE. (2) Further, we compare the training efficiency of the SASRec model using CCE and $\mathrm{CCE}^{-}$ over CE, $\mathrm{CE}^{-}$ in Sections 5.2. The results of this section further highlight the effectiveness of negative sampling in SRS. (3) Since uniform negative sampling is not the only possible sampling strategy in Section 5.4.1, we study other negative sampling strategies. (4) Finally, we extend our empirical analysis to the Bert4Rec model in Section 5.6.
5.1. Optimal Accuracy-Memory Trade-off
Figure 5. Memory scaling across three dimensions: (1) marker size (slsl: sequence length), (2) color (bsbs: batch size), and (3) marker shape (nsns: number of negative samples). Asterisks denote points using the CE loss. For example, brown markers (larger batch sizes) generally yield better average performance, though the largest batch size does not consistently dominate. Higher NDCG occurs with 128–512 negative samples (squares-hexagons); only the Zvuk dataset shows improved performance with asterisks (full $\mathrm{CE}^{-}$ loss).
Figure 6. This figure presents aggregated results from Figure 5, averaging across sequence lengths. Error bars indicate standard deviations. Colors represent batch sizes ($bs$). Performance increases with negative sample size ($ns$) for some datasets, while saturating for others (Movielens-20M, Gowalla). Batch size scaling shows no universal pattern: larger batches generally improve performance but saturate for 30Music, while smaller batches (128 vs 512) yield better results for Movielens-20M
The first research question is how to determine the optimal configuration of training hyperparameters given the memory constraints and how specific memory-related hyperparameters affect performance. We investigate how batch size (bs), the number of negatives (ns), and sequence length (sl) affect the NDCG@10 metric of the SASRec model after training. We used a fixed grid of configurations, evaluating approximately $170$ points per dataset, resulting in a total of $997$ runs across all datasets. While the grid does not provide a fully uniform coverage of the hyperparameter space, the analysis can still be interpreted under the assumption of reasonably even exploration (scarlett2017lower; srinivas2010gaussian).
Figure 5 demonstrates how bs, sl, and ns affect both training memory consumption and ranking quality (NDCG@10) across six datasets.
Points representing the use of the CE loss are indicated by asterisks on the plots. Analysis of all six datasets reveals several consistent and robust trends. A clear positive correlation exists between memory consumption and ranking quality: higher memory usage generally corresponds to improved NDCG@10 scores. This trend is consistently observed across all datasets.
Bs stands out as one of the most influential factors. Larger bs (512 and above), depicted in warm colors (red and orange dots), are generally linked to higher performance. This trend is especially evident in the Gowalla, Beauty, and Megamarket datasets, where red dots consistently appear near the top of the chart regardless of sl or ns. In contrast, smaller bs (e.g., 32 or 64), shown in cool blue tones, tend to cluster toward the lower part of the chart across almost all datasets, even when other parameters are increased.
Increasing the sequence length (512 and 1024) during training often leads to improved performance, particularly when combined with a batch size of 256 or greater. However, in most cases, the optimal approach involves using a large batch size with a long sequence, assuming sufficient resources are available.
Regarding the number of negative samples, the best results are typically achieved with moderate values (between 64 and 512). This is evident in datasets like Gowalla and Beauty, where the highest points correspond to shapes representing medium ns values. In contrast, extreme values, very small (e.g., 32) or very large numbers of negatives, tend to yield unstable and often weaker results. Finally, optimal configurations across all datasets are generally achieved by increasing bs and, in some cases, sl within the limits of available memory. Importantly, increasing the number of ns does not always result in improved performance. Low values across all three parameters consistently result in poor performance, while excessively increasing one without balancing the others often leads to inefficient memory usage and suboptimal NDCG@10 scores.
Based on the analysis, the following conclusion can be drawn.
General increase of NDCG@10 with growing memory consumption can be observed for Zvuk, MovieLens 20M, 30Music, and MegaMarket, though this increase is not linear. However, the plots for Beauty and Gowalla reveal saturation points, after which increasing resources provide diminishing returns.
Increasing batch size shares a similar trend. However, there is no guarantee across all considered datasets: on the Zvuk dataset, increasing bs beyond 128 offers minimal gains compared to 512, and for MovieLens 20M, the metric remains relatively stable.
Effect of negative sampling on the NDCG metric. For that, we conduct a separate analysis for the $\mathrm{CE}^{-}$ loss function. Figure 6 presents the relationship between ns and NDCG. Again, the impact of ns varies across datasets. In most cases, the metric increases with ns up to a certain point, after which it stabilizes or slightly decreases with a more prominent decrease for Gowalla, likely due to overfitting. For the MegaMarket dataset, the influence of ns is also less prominent. Minor variations in the metric values depending on sl suggest a low model sensitivity to this parameter.
Should we increase several dimensions simultaneously? To summarize specific dependencies, we use two approaches: Spearman correlation analysis and feature importance from the Random Forest model. The correlation analysis quantifies the relationships between training hyperparameters and evaluation metrics (NDCG@10, Coverage@10, and Surprisal@10). Complementarily, feature importance is assessed using the Random Forest model, which utilizes hyperparameters as input to predict quality metrics. The models were constructed separately for each dataset and achieved an $R^{2}>0.95$, highlighting the relative contribution of the hyperparameters in the success of a SRS.
Figures 7(a) and 7(b) present the results of the Spearman correlation analysis and feature importance as determined by the Random Forest model. The bs and the ns influence model quality the most, followed by sequence length. The importance of interactions such as bs-ns and sl-bs also stands out, indicating that optimizing these combinations can further enhance performance.
In summary, we recommend prioritizing adjustments to bs and ns, followed by calibrating sl and its interactions. For large item catalog datasets like Gowalla, 30Music, Zvuk, and Megamarket, increasing training hyperparameter values improves model accuracy. However, using CE or $\mathrm{CE}^{-}$ with large hyperparameter values incurs significant memory costs. Therefore, a memory-optimized approach is essential for efficient training on these datasets.
(a) Spearman correlations between parameters: sl, bs, ns and their pairwise products.
(b) Feature importance (dataset average) from Random Forest Regression predicting NDCG@10, Coverage@10, and Surprisal@10.
Figure 7. Parameter analysis: (a) correlation structure, (b) predictive importance for recommendation metrics, nsbs,nssl,slbsns-bs,ns-sl,sl-bs denote pairwise products as dependent variables.
Table 3. Comparison of CE and CCE. Hyperparameters such as $bs,sl$, and $ns$ were chosen based on our initial pool of scaling experiments for CE and $\mathrm{CE}^{-}$, respectively. CCE and $\mathrm{CCE}^{-}$ were trained using the same settings. We report NDCG@10, allocated memory (GB) and time per epoch (seconds). Bold values indicate best performance per dataset. $\Delta$: Relative difference compared to CE and $\mathrm{CE}^{-}$, green denotes CCE and $\mathrm{CCE}^{-}$ superiority.
Dataset Parameters CE CCE $\Delta$ NDCG, % $\Delta$ Mem, % $\Delta$ Time, %
(lr)3-5 (lr)6-8 (bs/sl) NDCG@10 Mem, GB Time, s NDCG@10 Mem, GB Time, s
Movielens-20m 256/512 0.0510 19.3 49.9 0.0505 2.5 37.0 0.0% ↓ 87.0% ↓ 25.9%
Beauty 1024/32 0.0114 51.7 132.0 0.0104 2.7 80.3 ↓ 8.6% ↓ 97.2% ↓ 39.2%
Gowalla 256/64 0.0225 47.7 57.5 0.0227 1.7 32.6 0.0% ↓ 96.4% ↓ 43.3%
Zvuk 64/96 0.0797 65.2 1766.1 0.0803 4.4 509.2 0.0% ↓ 93.3% ↓ 71.2%
30Music 32/128 0.0891 69.8 252.4 0.0841 4.4 118.8 ↓ 5.6% ↓ 93.7% ↓ 52.9%
Megamarket 16/128 0.0052 64.7 4054.3 0.0038 15.7 2936.8 ↓ 26.6% ↓ 75.7% ↓ 27.6%
Dataset Parameters $\mathrm{CE}^{-}$ $\mathrm{CCE}^{-}$ $\Delta$ NDCG, % $\Delta$ Mem, % $\Delta$ Time, %
(lr)3-5 (lr)6-8 (bs/sl/ns) NDCG@10 Mem, GB Time, s NDCG@10 Mem, GB Time, s
Movielens-20m 64/512/1511 0.0513 39.0 233.0 0.0488 1.6 87.0 ↓ 4.9% ↓ 95.9% ↓ 62.7%
Beauty 1024/32/511 0.0123 11.7 36.8 0.0117 2.7 23.0 ↓ 5.1% ↓ 87.2% ↓ 37.5%
Gowalla 512/256/511 0.0277 42.7 36.7 0.0276 5.3 36.0 0.0% ↓ 87.6% ↓ 23.4%
Zvuk 1024/32/255 0.0593 56.0 58.0 0.0566 6.9 35.0 ↓ 4.5% ↓ 87.7% ↓ 71.7%
30Music 256/128/511 0.1107 47.6 49.6 0.1132 3.6 18.6 ↑ 2.2% ↓ 92.4% ↓ 62.5%
Megamarket 1024/32/255 0.0117 71.0 60.0 0.0117 8.9 30.0 0.0% ↓ 87.5% ↓ 50%
Table 4. Comparison NDCG@10, memory consumption and epoch time for CCE and $\mathrm{CCE}^{-}$ on the SASRec model using identical training hyperparameters, except $ns$ for $\mathrm{CCE}^{-}$. Bold values indicate the best performance per dataset. $\Delta$: Relative difference compared to $\mathrm{CCE}^{-}$, green denotes $\mathrm{CCE}^{-}$ superiority.
Dataset Parameters CCE $\mathrm{CCE}^{-}$ $\Delta$ NDCG, $\Delta$ Mem, $\Delta$ Time,
(lr)3-5 (lr)6-8 (bs/sl/ns) NDCG Mem, Time, NDCG Mem, Time, % % %
@10 GB s @10 GB s
Movielens-20m 256/512/1511 0.0511 3.4 43.0 0.0523 3.5 52.0 ↑ 2.3% ↑ 2.9% ↑ 20.9%
Beauty 1024/32/511 0.0104 2.7 80.3 0.0116 2.7 23.0 ↑ 11.5% ↑ 7.1% ↓ 71.4%
Gowalla 1024/320/511 0.0241 11.8 132.0 0.0279 11.1 21.0 ↑ 15.8% ↑ 2.5% ↓ 84.1%
Zvuk 1024/128/2047 0.0940 10.6 550.0 0.0852 32.3 308.0 ↓ 10.6% ↑ 47.4% ↓ 44.0%
30Music 256/720/4095 0.1393 9.4 389.0 0.1436 58.4 277.0 ↑ 3.1% ↑ 261.7% ↓ 28.8%
Megamarket 5120/320/1023 0.0181 55.0 2248.0 0.0168 78.0 262.0 ↓ 7.2% ↑ 100.0% ↓ 88.3%
Table 5. Comparison Coverage@10 and Surprisal@10 time for CCE and $\mathrm{CCE}^{-}$ on the SASRec model using identical training hyperparameters, except $ns$ for $\mathrm{CCE}^{-}$. Bold values indicate the best performance per dataset. $\Delta$: Relative difference compared to $\mathrm{CCE}^{-}$, green denotes $\mathrm{CCE}^{-}$ superiority.
Dataset Parameters CCE $\mathrm{CCE}^{-}$ $\Delta$ Cov., $\Delta$ Surp.,
(lr)3-4 (lr)5-6 (lr)7-7 (lr)8-8 (bs/sl/ns) Coverage Surprisal Coverage Surprisal % %
@10 @10 @10 @10
Movielens-20m 256/512/1511 0.2102 0.0793 0.2429 0.0801 ↑ 15.6% ↑ 1.0%
Beauty 1024/32/511 0.1155 0.5172 0.1813 0.5069 ↑ 57.0% ↓ 2.0%
Gowalla 1024/320/511 0.2021 0.6447 0.2770 0.6276 ↑ 37.1% ↓ 2.7%
Zvuk 1024/128/2047 0.1401 0.3547 0.1490 0.3486 ↑ 6.3% ↓ 1.7%
30Music 256/720/4095 0.1203 0.5783 0.1241 0.5834 ↑ 3.2% 0%
Megamarket 5120/320/1023 0.1705 0.5466 0.1158 0.4941 ↓ 32.1% ↓ 9.6%
[…]
Although CE/CCE and $\mathrm{CE}^{-}$$\mathrm{CCE}^{-}$ are mathematically equivalent, their implementation code for GPU computation differs. This leads to variations in memory footprint, computational time, and numerical deviations, motivating a detailed comparison.
[…]
Memory consumption significantly contracts across all datasets for the CCE, indicating that the memory footprint is mainly due to the materialization of logits for computing the CE loss. The time per epoch also decreases by a minimum of 25.9%. This suggests that the fusion of the log-sum-exp operation with matrix multiplication and gradient filtering implemented in the CCE kernels enhances the efficiency of computing the CE loss.
Significant reduction of computational resources occurs for $\mathrm{CCE}^{-}$ compared to $\mathrm{CE}^{-}$. The memory footprint decreases by no less than 87% for all datasets. The time per epoch also reduces by more than 23%, achieving 73% for Zvuk.
Deviation in metrics: Although resource savings are evident, NDCG@10 performance slightly decreases on some datasets when training with CCE or $\mathrm{CCE}^{-}$ compared to CE and $\mathrm{CE}^{-}$. The only exception is the 30Music dataset, where $\mathrm{CCE}^{-}$ shows an improvement over $\mathrm{CE}^{-}$.
Since the loss function formulation in CCE and $\mathrm{CCE}^{-}$ is simply a logarithmic decomposition of the $\mathrm{LogSoftmax}$ function used in the PyTorch implementation of CE and $\mathrm{CE}^{-}$, the observed metrics discrepancies may arise from minor numerical inaccuracies during intermediate computations in Triton kernels or from loss scaling in mixed-precision training. However, this issue requires further investigation, which will be addressed in future work.
Table 6. Results for $\mathrm{CCE}^{-}$ with popularity sampling on the SasRec model using identical hyperparameters as for global uniform sampling
Dataset Parameters (bs/sl/ns) NDCG@10 Coverage@10 Surprisal@10 Mem (GB) Time (s)
Movielens-20m 256 / 512 / 1511 0.0235 0.2221 0.1337 6.2 111.0
Beauty 1024 / 32 / 511 0.0066 0.5181 0.7098 5.7 82.9
Gowalla 1024 / 320 / 511 0.0200 0.4540 0.7521 67.9 86.9
Table 7. Results for the comparison of $\mathrm{CCE}^{-}$ loss with global uniform and popularity sampling for the SASRec model. $\Delta$: Relative difference compared to popularity sampling and global uniform sampling, green denotes global uniform sampling superiority.
Dataset (bs/sl) $\Delta$ NDCG, % $\Delta$ Coverage, % $\Delta$ Surprisal, % $\Delta$ Mem, % $\Delta$ Time, %
Movielens 20m 256 / 512 ↑ 55.1% ↑ 8.6% ↓ 66.9% ↓ 77.1% ↓ 113.5%
Beauty 1024 / 32 ↑ 43.1% ↓ 185.8% ↓ 40.0% ↓ 111.1% ↓ 260.4%
Gowalla 1024 / 320 ↑ 28.3% ↓ 63.9% ↓ 19.8% ↓ 511.7% ↓ 313.8%
5.3. Reallocating memory savings to increase performance with CCE and $\mathrm{CCE}^{-}$
The results discussed in the previous section indicate that CCE and $\mathrm{CCE}^{-}$ significantly reduce the memory consumption required for training SASRec. This reduction provides an opportunity to enhance model performance by increasing batch size ($bs$), sequence length ($sl$), and number of samples ($ns$) in the case of $\mathrm{CCE}^{-}$, implementing insights from the Section 5.1. In these experiments, we set identical training hyperparameters for both CCE and $\mathrm{CCE}^{-}$. Tables 4, 5 present the values of the NDSG@10, Coverage@10, and Surprisal@10 metrics for all datasets, along with the allocated memory and time per epoch required for training.
Improvement in model performance across all datasets is observed when comparing results in Tables 3,  4, 5 and in Figure 1, except for smaller Beauty, where adjustments to the training parameters do not yield an increase in metrics. The most significant improvements are observed in Megamarket, which shows a 54.7% increase, followed by 30Music with a 27% increase, and Zvuk with a 17.1% increase. These datasets are characterized by extensive catalogs, exceeding 800K items. Conversely, for datasets Beauty, Gowalla, ML-20M, which have item catalogs not exceeding 310K items, applying CCE and $\mathrm{CCE}^{-}$ for model training does not lead to considerable improvement in their performance. Thus, using the CCE and $\mathrm{CCE}^{-}$ for training models with large final layers improves performance, while for smaller catalogs main benefit is training acceleration.
Negative sampling during training with the $\mathrm{CCE}^{-}$ accelerates training compared to CCE across datasets except Movielens-20m (ML20M), which includes only 15K items. For example, using 1023 negative samples for Megamarket results in a time per epoch reduction of 88.3% compared to CCE, while using 4095 negative samples for 30Music leads to only a 28.8% acceleration. Additionally, restricting the item catalog during scoring may degrade model performance, particularly for datasets with large item catalogs like Megamarket and Zvuk. In contrast, for datasets with fewer items, such as 30Music, Gowalla, and Beauty, training with $\mathrm{CCE}^{-}$ can yield improvements in model metrics compared to CCE.
What is better $\mathrm{CCE}^{-}$ over CCE? We show that $\mathrm{CCE}^{-}$ is faster and outperforms CCE on four out of six datasets due to negative sampling, however, it requires more memory to store a matrix of negative sample indices compared to CCE.
5.4. Gradient Filtering with CCE
Here we attempt to answer Q2 posed in Section 1. To quantify the impact of pruning near-zero gradients, we conducted experiments using $\mathrm{CCE}^{-}$ across six benchmark datasets. The training hyperparameters were kept consistent with those used in the CCE and $\mathrm{CCE}^{-}$ comparison experiments. To analyze gradient sparsity, we performed a series of experiments where the threshold for zeroing out gradients was incrementally increased in each run.
Figure 8 shows training time per epoch and NDCG@10 metric of the model. We can observe two trends: (1) Performance stability: Model accuracy degrades only when the threshold exceeds $1\times 10^{-4}-1\times 10^{-3}$. (2) Efficiency gains: The training time per epoch consistently enhances as computations of near-zero gradients are omitted.
These results suggest that gradients close to zero contribute minimally to model updates during later training stages, enabling safe pruning for computational savings. Moreover, for Movielens-20M gradient pruning induces regularization that leads to improved NDCG@10 metrics when the pruning threshold increases from $1\times 10^{-6}-1$ to $1\times 10^{-4}-1$. This indicates that for this dataset, there are many noisy gradients that complicate training.
Figure 8. Performance of SASRec trained with CCE using gradient filtering at increasing thresholds (filter_eps parameter). The NDCG@10 metric (left axis) and the training time per epoch (right axis) are shown.
Since CCE with gradient pruning enables accelerating the training process, we compare the time per epoch and the accuracy of models trained with CCE and $\mathrm{CCE}^{-}$. The training hyperparameters matched those specified in Table 4.
Based on the results shown in Figure9, it can be concluded that $\mathrm{CCE}^{-}$ achieves NDCG@10 scores comparable to those of CCE. At that training with $\mathrm{CCE}^{-}$ requires the least time per epoch on five of the six datasets, with the exception of Movielens-20m. For the Beauty, Megamarket, and Gowalla datasets, $\mathrm{CCE}^{-}$ achieves over 3x acceleration over CCE despite of using gradient filtering at a threshold of $1\times 10^{-4}$. Under these conditions, the NDCG@10 score for $\mathrm{CCE}^{-}$ is higher across all three datasets, with the smallest observed improvement, an 8% increase, occurring in the Megamarket dataset.
For the Movielens-20m dataset, training with CCE is approximately 35% faster per epoch than $\mathrm{CCE}^{-}$, even without gradient filtering. This dataset features the smallest item catalog among the considered datasets, comprising approximately 15K items 2. At that, $\mathrm{CCE}^{-}$ requires $1511$ negative samples to reach an NDCG@10 score comparable to that of $\mathrm{CCE}^{-}$. The large number of negative samples relative to the catalog size causes significant overhead when computing the forward and backward passes in $\mathrm{CCE}^{-}$ for the classification layer.
Figure 9. NDCG@10 metric and time per epoch of the SASRec models trained with $\mathrm{CCE}^{-}$ and CCE using gradient filtering at increasing thresholds.
5.4.1. Comparison Sampling Strategies
One of the main features of training with the CE loss using negative sampling is the ability to employ diverse strategies for selecting negative samples. In this work, we primarily focus on the global uniform strategy, where negative samples are drawn uniformly across the entire dataset.
As an alternative, we also consider popularity-based negative sampling, where negative samples are selected based on item popularity. In this case, the probability of selection increases with an item’s popularity, causing popular items to be chosen as negative samples more frequently.
Tables 6, 7 present data to compare metrics and computational resource usage for training with $\mathrm{CCE}^{-}$ using global uniform and popularity-based sampling strategies. In terms of $\mathrm{NDSG@10}$ metric, global uniform sampling consistently outperforms the popularity-based approach across all considered datasets. However, popularity-based sampling shows significant gains in diversity-related metrics for the Gowalla and Beauty datasets. Compared to popularity-based sampling, using global uniform sampling leads to a reduction in $\mathrm{Coverage@10}$ by 185.8% and 63.9 %, and a reduction in $\mathrm{Surprisal@10}$ by 40% and 19.8%. These results align with observations that popularity-weighted sampling enhances exposure to long-tail items, improving recommendation diversity and novelty (wang2024popularitylearning).
The popularity sampling strategy demands higher memory consumption during training compared to the global uniform sampling strategy. This increased requirement is due to the implementation characteristics of torch.multinomial, which necessitates the materialization of a tensor representing the unnormalized probabilities for sampling negative items in popularity-based scenarios. Consequently, only datasets with relatively small item catalogs can be processed under this strategy.
Another approach to negative sampling is the Scalable Cross-Entropy (SCE) method  (mezentsev2024scalable). This method approximates the CE loss by employing a sampling procedure that computes logits only for elements with the largest absolute values, rather than for all items in the catalog. As a result, memory usage for the logit tensor is significantly reduced, and the loss computation is accelerated.
Table 8 presents the results of the SASRec model trained with the SCE loss. Tables 9, 10 list the relative differences calculated for the metrics achieved by CCE and $\mathrm{CCE}^{-}$.
When comparing SCE with $\mathrm{CCE}^{-}$, $\mathrm{CCE}^{-}$ achieves considerably higher $\mathrm{NDCG@10}$ scores on four out of six datasets, with the exception of Zvuk and Megamarket. However, SCE provides better $\mathrm{Suprisal@10}$ metrics for all datasets.
Compared to $\mathrm{CCE}^{-}$, SCE requires more memory on the Megamarket dataset. Therefore, training with SCE was conducted using a batch size of 2048. In this case, SCE achieves $\mathrm{NDSG@10}$ score comparable with $\mathrm{CCE}^{-}$, which is applied for batch size 5120.
For CCE $\mathrm{NDSG@10}$ values are higher than those of SCE across all datasets. On the Megamarket CCE with batch size 5120 also outperforms SCE on 11.6 % in terms of $\mathrm{NDSG@10}$. Moreover, CCE requires significantly less memory during training, reducing memory usage by approximately 80% on both the 30Music and Beauty datasets.
The time per epoch for SCE is considerably lower than for CCE and $\mathrm{CCE}^{-}$ on datasets with large item catalogs, such as Zvuk, 30Music, and Megamarket. Therefore, this sampling approach can offer improved computational efficiency when training RecSys models with a large number of logits. Since SCE incorporates the CE loss function, it could be further optimized in the future using our Triton kernels, which compute the CE loss with negative sampling by using only the indices of positive and negative samples. This eliminates the need to materialize logits, thereby reducing memory consumption during training.
Table 8. Results for the SCE loss with the SASRec model
Dataset Parameters SCE
(lr)3-7 (bs/sl) NDCG@10 Coverage@10 Surprisal@10 Mem, GB Time, s
Megamarket 1024 / 32 0.0124 0.0706 0.5278 14.5 36.4
Megamarket 2048 / 320 0.0160 0.0988 0.5358 57.0 144.2
Zvuk 1024 / 32 0.0781 0.0804 0.3725 10.1 34.0
Zvuk 1024 / 128 0.0879 0.1109 0.3711 13.5 68.4
30Music 256 / 128 0.1011 0.0912 0.5682 8.1 13.7
30Music 256 / 720 0.1218 0.1065 0.5812 16.7 45.5
Gowalla 512 / 256 0.0205 0.2805 0.6676 7.9 14.8
Gowalla 1024 / 320 0.0222 0.3463 0.6887 15.9 23.6
Beauty 1024 / 32 0.0086 0.2290 0.5746 4.9 167.8
Movielens 20m 64 / 512 0.0477 0.1262 0.0670 3.6 34.2
Movielens 20m 256 / 512 0.0396 0.1812 0.0856 5.4 69.5
Table 9. Results for the comparison of SCE loss with $\mathrm{CCE}^{-}$ for the SASRec model. $\Delta$: Relative difference compared to SCE and $\mathrm{CCE}^{-}$, green denotes $\mathrm{CCE}^{-}$ superiority.
Dataset (bs/sl) $\Delta$ NDCG, % $\Delta$ Coverage, % $\Delta$ Surprisal, % $\Delta$ Mem, % $\Delta$ Time, %
Movielens 20m 256 / 512 ↑ 24.3% ↑ 25.9% ↓ 6.9% ↓ 54.3% ↓ 33.7%
Beauty 1024 / 32 ↑ 25.9% ↓ 26.3% ↓ 13.4% ↓ 81.4% ↓ 629.5%
Gowalla 1024 / 320 ↑ 20.4% ↓ 25.1% ↓ 9.7% ↓ 43.2% ↓ 12.4%
Zvuk 1024 / 128 ↓ 3.2% ↑ 25.6% ↓ 6.5% ↑ 58.2% ↑ 77.8%
30Music 256 / 128 ↑ 15.2% ↑ 14.2% 0.0% ↑ 71.4% ↑ 83.6%
[…]
Table 10. Results for the comparison of SCE loss with CCE for the SASRec model. $\Delta$: Relative difference compared to SCE and $\mathrm{CCE}^{-}$, green denotes CCE superiority.
Dataset (bs/sl) $\Delta$ NDCG, % $\Delta$ Coverage, % $\Delta$ Surprisal, % $\Delta$ Mem, % $\Delta$ Time, %
Movielens 20m 256 / 512 ↑ 22.5% ↑ 13.8% ↓ 7.9% ↓ 58.8% ↓ 61.6%
Beauty 1024 / 32 ↑ 17.3% ↓ 98.3% ↓ 11.1% ↓ 81.5% ↓ 109.0%
Gowalla 1024 / 320 ↑ 7.9% ↓ 71.4% ↓ 6.8% ↓ 34.7% ↑ 82.1%
Zvuk 1024 / 128 ↑ 6.5% ↑ 20.8% ↓ 4.6% ↓ 27.4% ↑ 87.6%
30Music 256 / 720 ↑ 12.6% ↑ 11.5% ↓ 0.5% ↓ 77.7% ↑ 88.3%
[…]
5.5. Comparison of CCE and $\mathrm{CCE}^{-}$ with BCE
As outlined in the original paper (kang2018self), the SASRec model can be trained using the Binary Cross-Entropy (BCE) loss. To evaluate the effectiveness of our proposed methods, we conducted training SASRec with the BCE loss.
Table 11 summarizes the performance of SASRec trained with the BCE loss. A comparison with models trained using the $\mathrm{CCE}^{-}$ and CCE losses is presented in Tables 12 and 13. Under identical settings for batch size and sequence length, SASRec with the BCE loss achieves significantly lower NDCG@10 and Coverage@10 scores compared to both $\mathrm{CCE}^{-}$ and CCE for all datasets.
This performance gap is particularly pronounced on the Megamarket, Zvuk, and Beauty datasets. Notably, on Beauty, using $\mathrm{CCE}^{-}$ or CCE leads to improvements in NDCG@10 by 215.5% and 182.8%, respectively, and in Coverage@10 by 4933% and 3106%.
The Surprisal@10 metric is also lower for all datasets, with the exception of the Movielens-20m dataset. For this dataset, the metrics are higher by 8.1% relative to $\mathrm{CCE}^{-}$ and by 9.0% relative to CCE.
However, since BCE uses only one negative sample, training with this loss function requires considerably less time per epoch than $\mathrm{CCE}^{-}$ and CCE. Higher slowdown is 1182% for the 30Music dataset. For Megamarket, a usage of CCE leads to a rise in training epoch by 3935.9% compared with BCE. Notably, memory consumption reduces by 8.5%. Memory consumption for CCE is higher only for training on 30Music, Gowalla, Beauty. On the other hand, $\mathrm{CCE}^{-}$ requires more memory than BCE for all datasets with the exception Movielens-20M.
However, since the BCE loss uses only a single negative sample per positive instance, training with this loss is significantly faster per epoch compared to $\mathrm{CCE}^{-}$ and CCE. The highest observed slowdown for $\mathrm{CCE}^{-}$ is 1182% on the 30Music dataset. Conversely, for the Megamarket dataset, using CCE increases the training time per epoch by 3935.9% compared to BCE. Despite this slowdown, using CCE leads to only an 8.5% decrease in memory efficiency.
In fact, CCE requires more memory than BCE only on the 30Music, Gowalla, and Beauty datasets. In contrast, $\mathrm{CCE}^{-}$ consistently demands more memory across all datasets except Movielens-20M.
Table 11. Results for the BCE loss with the SASRec model.
Dataset Parameters BCE
(lr)3-7 (bs/sl) NDCG@10 Coverage@10 Surprisal@10 Mem, GB Time, s
Megamarket 1024 / 32 0.004294 0.049178 0.376738 14.3 26.9
Megamarket 5120 / 320 0.007127 0.055880 0.378771 60.1 55.7
Zvuk 1024 / 32 0.025003 0.055162 0.299418 9.2 34.8
Zvuk 1024 / 128 0.030669 0.078245 0.336043 11.2 31.8
30Music 256 / 128 0.080305 0.058708 0.480891 6.5 8.9
30Music 256 / 720 0.091240 0.067944 0.498069 8.8 21.6
Gowalla 512 / 256 0.017491 0.121836 0.569349 4.7 14.2
Gowalla 1024 / 320 0.019765 0.127281 0.568624 9.5 14.3
Beauty 1024 / 32 0.003677 0.003602 0.452931 2.4 30.4
Movielens 20m 64 / 512 0.029959 0.189254 0.087503 1.5 66.9
Movielens 20m 256 / 512 0.032907 0.190266 0.087117 3.5 38.3
Table 12. Results for the comparison of BCE loss with $\mathrm{CCE}^{-}$ for the SASRec model. $\Delta$: Relative difference compared to BCE and $\mathrm{CCE}^{-}$, green denotes $\mathrm{CCE}^{-}$ superiority.
(bs/sl) $\Delta$ NDCG, % $\Delta$ Coverage., % $\Delta$ Surprisal., % $\Delta$ Mem, % $\Delta$ Time, %
Megamarket 5120 / 320 ↑ 135.7% ↑ 107.2% ↑ 30.4% ↑ 29.8% ↑ 370.4%
Zvuk 1024 / 128 ↑ 177.8% ↑ 90.4% ↑ 3.7% ↑ 188.4% ↑ 868.6%
30Music 256 / 720 ↑ 57.4% ↑ 82.7% ↑ 17.1% ↑ 563.6% ↑ 1182.4%
Gowalla 1024 / 320 ↑ 41.2% ↑ 117.6% ↑ 10.4% ↑ 16.8% ↑ 46.9%
Beauty 1024 / 32 ↑ 215.5% ↑ 4933.3% ↑ 11.9% ↑ 12.5% ↓ 24.3%
Movielens 20m 256 / 512 ↑ 58.9% ↑ 27.7% ↓ 8.1% 0.0% ↑ 35.8%
Table 13. Results for the comparison of BCE loss with CCE for the SASRec model. $\Delta$: Relative difference compared to BCE and CCE, green denotes CCE superiority.
(bs/sl) $\Delta$ NDCG, % $\Delta$ Coverage., % $\Delta$ Surprisal., % $\Delta$ Mem, % $\Delta$ Time, %
Megamarket 5120 / 320 ↑ 154.0% ↑ 205.1% ↑ 44.3% ↓ 8.5% ↑ 3935.9%
Zvuk 1024 / 128 ↑ 206.5% ↑ 79.1% ↑ 5.6% ↓ 5.4% ↑ 1629.6%
30Music 256 / 720 ↑ 52.7% ↑ 77.1% ↑ 16.1% ↑ 6.8% ↑ 1700.9%
Gowalla 1024 / 320 ↑ 21.9% ↑ 58.8% ↑ 13.4% ↑ 24.2% ↑ 823.1%
Beauty 1024 / 32 ↑ 182.8% ↑ 3106.6% ↑ 14.2% ↑ 12.5% ↑ 164.1%
Movielens 20m 256 / 512 ↑ 55.3% ↑ 10.5% ↓ 9.0% ↓ 2.9% ↑ 12.3%
5.6. Analysis of efficiency of CCE and $\mathrm{CCE}^{-}$ for Bert4Rec
In addition to the SASRec architecture, we study the efficiency of CCE and $\mathrm{CCE}^{-}$ for Bert4Rec  (sun2019bert4rec).
Table 14 presents a comparative analysis of CCE and $\mathrm{CCE}^{-}$ on the Bert4Rec model using an identical batch size and sequence length. It is evident that $\mathrm{CCE}^{-}$ allows for a reduction in training time on most datasets due to negative sampling. For example, on Megamarket, there is a 3.8% improvement in quality $\mathrm{NDCG@10}$ alongside a 14.4% decrease in memory consumption and a 32.7% reduction in training time. The quality gain is especially pronounced on Gowalla (68.5%). At the same time, some datasets (Zvuk, 30Music) show a slight drop in quality metrics with a significant time gain.
For completeness, we also provide metrics for Bert4Rec trained with CE, $\mathrm{CE}^{-}$ in Tables 16, 17. Table 16 indicates that using the full CE achieves quality comparable to CCE on datasets with relatively small item catalogs, such as Movielens-20m, Beauty, and Gowalla. On datasets with large item catalogs, the performance of CE is considerably worse than that of CCE, which enables the use of significantly larger batch sizes and sequence lengths.
According to Table 17, $\mathrm{CE}^{-}$ exhibits nearly zero NDCG@10 values on most datasets, indicating that this loss is unsuitable for the task. In contrast, $\mathrm{CCE}^{-}$ shows results comparable to CE. This may be due to batch size and the number of negatives: smaller batch sizes and fewer negatives lead to a strong drop in $\mathrm{NDCG@10}$ quality.
The results demonstrate that CCE and $\mathrm{CCE}^{-}$ are effective for Bert4Rec. These approaches improve the model’s predictive quality on datasets with large catalogs by enabling the use of larger batch sizes and longer sequence lengths through memory optimization.
Table 14. Comparison NDCG@10, memory consumption and epoch time for CCE and $\mathrm{CCE}^{-}$ on the Bert4Rec model using identical training hyperparameters. Bold values indicate best performance per dataset. $\Delta$: Relative difference compared to $\mathrm{CCE}^{-}$, green denotes $\mathrm{CCE}^{-}$ superiority.
Dataset Parameters CCE $\mathrm{CCE}^{-}$ $\Delta$ NDCG, $\Delta$ Mem, $\Delta$ Time,
(lr)3-5 (lr)6-8 (bs/sl/ns) NDCG Mem, Time, NDCG Mem, Time, % % %
@10 GB s @10 GB s
Mega Market 2048 / 320 / 4095 0.0125 62.8 306.3 0.0130 53.8 206.1 ↑ 3.8% ↓ 14.4% ↓ 32.7%
Zvuk 1024 / 128 / 2047 0.0793 10.2 101.6 0.0639 10.6 65.8 ↓ 19.5% ↑ 3.9% ↓ 35.2%
30Music 256 / 720 / 4095 0.0885 11.6 55.9 0.0803 12.3 51.1 ↓ 9.2% ↑ 6.0% ↓ 8.6%
Gowalla 1024 / 320 / 511 0.0092 10.7 13.0 0.0155 10.7 10.6 ↑ 68.5% 0% ↓ 18.5%
Beauty 1024 / 32 / 511 0.0080 2.1 21.0 0.0085 2.1 21.3 ↑ 6.3% 0% ↑ 1.4%
Movielens-20m 256 / 512 / 511 0.0489 3.5 24.3 0.0481 3.5 26.3 ↓ 1.5% 0% ↑ 8.2%
Table 15. Comparison Coverage@10 and Surprisal@10 for CCE and $\mathrm{CCE}^{-}$ on the Bert4Rec model using identical hyperparameters. Bold values indicate best performance per dataset. $\Delta$: Relative difference compared to $\mathrm{CCE}^{-}$, green denotes $\mathrm{CCE}^{-}$ superiority.
Dataset Parameters CCE $\mathrm{CCE}^{-}$ $\Delta$ Cov., $\Delta$ Surp.,
(lr)3-4 (lr)5-6 (bs/sl/ns) Coverage Surprisal Coverage Surprisal % %
@10 @10 @10 @10
Mega Market 2048 / 320 / 4095 0.1410 0.5603 0.1352 0.5392 ↓ 4.1% ↓ 3.8%
Zvuk 1024 / 128 / 2047 0.1042 0.3578 0.1014 0.3435 ↓ 2.7% ↓ 4.0%
30Music 256 / 720 / 4095 0.0799 0.5726 0.0849 0.5600 ↑ 6.3% ↓ 2.2%
Gowalla 1024 / 320 / 511 0.1710 0.6548 0.1557 0.6109 ↓ 9.0% ↓ 6.7%
Beauty 1024 / 32 / 511 0.0502 0.5093 0.0301 0.4796 ↓ 40.0% ↓ 5.8%
Movielens-20m 256 / 512 / 511 0.1690 0.0744 0.1617 0.0724 ↓ 4.3% ↓ 2.8%
Table 16. Performance of Bert4Rec with CE loss.
Dataset Parameters CE
(lr)3-7 (bs/sl) NDCG@10 Coverage@10 Surprisal@10 Mem, GB Time, s
Megamarket 32 / 128 0.0035 0.0044 0.3603 62.7 3684.4
Zvuk 64 / 96 0.0507 0.0221 0.2644 38.1 1514.6
30Music 32 / 128 0.0237 0.0087 0.4026 15.4 183.2
Gowalla 256 / 64 0.0091 0.0953 0.6321 12.9 30.0
Beauty 1024 / 32 0.0083 0.0316 0.4875 12.6 62.8
Movielens-20m 256 / 512 0.0487 0.1890 0.0771 7.0 32.9
Table 17. Performance of Bert4Rec with $\mathrm{CE}^{-}$ loss.
Dataset Parameters $\mathrm{CE}^{-}$
(lr)3-7 (bs/sl) NDCG@10 Coverage@10 Surprisal@10 Mem, GB Time, s
Megamarket 1024 / 32 / 255 0.0040 0.0415 0.4278 35.4 67.1
Zvuk 1024 / 32 / 256 0.0001 0.0006 0.4146 22.8 86.4
30Music 256 / 128 / 511 0.0000 0.0000 0.6132 23.4 40.5
Gowalla 512 / 256 / 512 0.0001 0.0009 0.6236 9.9 13.2
Beauty 1024 / 32 / 512 0.0000 0.0009 0.7653 2.8 21.2
Movielens-20m 64 / 512 / 1512 0.0002 0.0257 0.1928 1.0 35.6
6. Conclusions
We proposed CCE and $\mathrm{CCE}^{-}$ for memory-efficient SRS training as a replacement for full CE loss and its variant with negative sampling correspondingly. By employing efficient Triton kernel implementations, these methods eliminate the memory bottleneck caused by materializing logit tensors. Our experiments show that CCE and $\mathrm{CCE}^{-}$ reduce the training memory footprint by 75–97% and accelerate training by 23–71% if compared to the conventional PyTorch implementation. Notably, $\mathrm{CCE}^{-}$’s ability to reduce the number of negative samples can further decrease epoch duration and enhance model performance compared to CCE, and CCE can be further optimized through gradient pruning in the final layer, providing a complementary strategy to minimize computational overhead.
Future research could focus on enhancing the computational efficiency of Triton kernels, integrating alternative negative sampling strategies, and developing adaptive gradient pruning techniques to balance sparsity and model performance.
Our experiments with SASRec and BERT4Rec demonstrate that memory optimization during training using CE-type losses can enhance sequential recommender systems. The insights and optimizations we employ, such as sparse gradients in CCE and the design of $\mathrm{CCE}^{-}$, are model-agnostic and can be seamlessly integrated into other architectures, ensuring the broader applicability of our findings to SRS.
7. Acknowledgements
The work was supported by the grant for research centers in the field of AI provided by the Ministry of Economic Development of the Russian Federation in accordance with the agreement  [^00] C313925P4F0002 and the agreement with Skoltech №139-10-2025-033.